OpenStudy (anonymous):

Two skaters race in different directions. The first skater travels at 18 mph and the second travels at 23 mph. How long will it be before they are 123 miles apart? a. 4 hours b. 3 hours c. 1 hour d. 2 hours

OpenStudy (anonymous):

b. 3 hours

OpenStudy (cwrw238):

speed at which they are travelling apart = 18 + 23 = 41 mph so time to travel 123 miles = 123/41 = 3 hours

OpenStudy (anonymous):

Thank you! The way I was doing it, I was totally wrong.

OpenStudy (cwrw238):

yw